A couple purchasing a home budget $1,200 per month for their loan payment. If they have $17,000 available for a down payment and are considering a 25-year loan, how much can they spend on the home at each of the f

(a)    6.6% compounded monthly

(b)    7.8% compounded monthly
Answer #1

a.

Calculating Present Value of Payment,

Using TVM calculation,

PV = [FV = 0, PMT = 1,200, N = 300, I 0.066/12]

PV = $176,090.16

Loan Amount = 17,000 + 176,090.16 = $193,090.16

b.

Calculating Present Value of Payment,

Using TVM calculation,

PV = [FV = 0, PMT = 1,200, N = 300, I 0.078/12]

PV = $158,183.14

Loan Amount = 17,000 + 158,183.14 = $175,183.14